OpenStudy (anonymous):

If you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the solution to the system? The solution is (4, 4) The solution is x = 4 and y = 4 There is no solution There are infinitely many solutions

OpenStudy (agl202):

If you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the system Sol: This means that the conditions are satisfied or point lies on the equation............ans Part 2: Solve the following system and show all of your work. 3y − 2x = 11...............(i) y + 2x = 9..................(ii) by adding 4y =20 or y = 5................Ans 3(5) -2x =11 -2x = 11-15 = -4 x = 2...............................Ans

OpenStudy (anonymous):

So does that mean that the answer is, there are infinitely many solutions?

OpenStudy (agl202):

yes!

OpenStudy (anonymous):

Are you there

OpenStudy (anonymous):

So it is infinitely many solutions ?

OpenStudy (agl202):

ya

OpenStudy (agl202):

Yes, thats correct!

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

Can I tag you in another question to help me with ?

OpenStudy (agl202):

sure